# Tight PAC-Bayesian Risk Certificates for Contrastive Learning

### v2

@@ 3.1.2 Second PAC-Bayes certificate
While theorem 3 uses the bounded difference assumption, an alternative approach would be to replace $\sum_{x^{\prime}\in X}\operatorname{sim}(x_{i},x^{\prime})$ in (1) by its expected value, conditioned on $x_{i}$, thereby making the terms in (1) independent. To implement this in the PAC-Bayes framework, we define an $\varepsilon$-modified SimCLR empirical loss $L_{S}^{\prime}(Q)=\int_{\mathcal{W}}L_{S}^{\prime}(f_{w})\,Q(dw)$ by replacing $\ell_{\text{cont}}$ in (1) with an upper bound $\ell^{\prime}_{\text{cont}}$ as
@@
where $B:=\frac{1}{\tau}+\log\left(me^{\frac{1}{\tau}}\right)+\varepsilon$ for $\varepsilon=\left(e^{\frac{1}{\tau}}-e^{-\frac{1}{\tau}}\right)\sqrt{\frac{m-1}{2\alpha}\log\frac{2}{\delta}}$
@@
Observe that Theorem 4 still holds for the original SimCLR loss by replacing $\varepsilon$ with $2\varepsilon$. This result follows from grouping terms involving augmented views of the same sample into one variable, $\operatorname{sim}(x,x^{\prime})+\operatorname{sim}(x,{x^{\prime}}^{+})$, before applying the concentration bound. This ensures that the variables are independent while multiplying the bound by $2$ and explains the factor $2$ in the $\varepsilon$ term.
The remainder of this subsection is dedicated to the proof of theorem 4. First, we apply a Hoeffding’s inequality to the term involving the negative samples, as stated below.
@@
Let $f\sim Q$ and recall the notation $S(x,X):=\sum\limits_{x^{\prime}\in X^{-}}\operatorname{sim}(x,x^{\prime})$ For all $\delta\in(0,1)$, with $\varepsilon=\left(e^{\frac{1}{\tau}}-e^{-\frac{1}{\tau}}\right)\sqrt{\frac{m-1}{2}\log\frac{1}{\delta}}$, the following concentration bound holds: $\mathbb{P}\left(S(x,X)-\mathbb{E}[S(x,X)]\geq\varepsilon\mid x\right)\leq\delta.$
@@
Recall the SimCLR population loss $L(f)$ defined in eq. 2. Define the intermediate loss $\widetilde{L}(f)$ which is a mean of i.i.d. terms obtained by replacing $\ell_{\text{cont}}$ with the loss ~$\tilde{\ell}_{\text{cont}}$ given by ~$\tilde{\ell}_{\text{cont}}(x,x^{+}):=-\log\frac{\operatorname{sim}(x,x^{+})}{\operatorname{sim}(x,x^{+})+\mathbb{E}\left[S(x,X)\mid x\right]+\varepsilon}$ for a suitable $\varepsilon>0$. We obtain the following
@@
Using lemma 5, the following lemma shows that $L(f)$ can be upper-bounded by $\widetilde{L}(f)$.
@@
Let $f\sim Q$ and $\widetilde{L}(f)$ defined in eq. 6. For $\delta\in(0,1)$ and $B_{\ell}=\frac{2}{\tau}+\log m$, we have
@@
For $\delta\in(0,1)$ and $B_{\ell}=\frac{2}{\tau}+\log m$,. Recall that $\ell_{\operatorname{cont}}(x,x^{+},X^{-})$ can be written as an average of two terms that play a symmetric role. Hence, we will first upper-bound $\ell_{\operatorname{cont}}(x,x^{+},X)$ where $X=\bigcup_{x^{\prime}\neq x}\{x^{\prime}\}$ Observe that $\ell_{\operatorname{cont}}(x,x^{+},X)$ can be split into two terms using $\varepsilon$ defined in lemma 5:
$\ell_{\operatorname{cont}}(x,x^{+},X)=\tilde{\ell}_{\text{cont}}(x,x^{+})\mathbb{I}_{\{S(x,X)<\mu(x)+\varepsilon\}}+\ell_{\operatorname{cont}}(x,x^{+},X)\mathbb{I}_{\{S(x,X)\geq\mu(x)+\varepsilon\}},$
where $\mu(x)=\mathbb{E}\left[S(x,X)\mid x\right]$.
Noting that $\ell_{\operatorname{cont}}(x,x^{+},X)\leq B_{\ell}$ and $\mathbb{I}_{\{\cdot\}}\leq 1$, we obtain $\ell_{\operatorname{cont}}(x,x^{+},X)\leq\tilde{\ell}_{\text{cont}}(x,x^{+})+B_{\ell}\mathbb{I}_{\{S(x,X)\geq\mu(x)+\varepsilon\}}$. We derive:
@@
where $X^{\prime}=\bigcup_{x^{\prime}\neq x}\{{x^{\prime}}^{+}\}$. Using the inequality established in lemma 5 and the expression of $L(f)$ combined with eq. 8, we have
$L(f)\leq\underset{(x_{i},x_{i}^{+})_{i=1}^{m}\sim\mathcal{S}^{m}}{\mathbb{E}}\left[\frac{1}{m}\sum_{i=1}^{m}\tilde{\ell}(x_{i},x_{i}^{+})\right]+B_{\ell}\delta$ (9)
@@
Observing that $\widetilde{L}_{S}(f)$ is an average of i.i.d. terms, the following result is a direct application of the PAC-Bayes-kl bound (theorem 1) after suitable rescaling of the loss function to the interval $[0,1]$.
@@
For a given dataset $S\sim\mathcal{S}^{n}$, let $\widetilde{L}(f)$ and $\widetilde{L}_{S}(f)$ be defined as in eq. 6 and eq. 10 respectively. We extend the definition of these losses to the PAC-Bayes framework: $\widetilde{L}(Q)=\mathbb{E}_{f\sim\mathcal{Q}}\left[\widetilde{L}(f)\right]$ and $\widetilde{L}_{S}(Q)=\mathbb{E}_{f\sim\mathcal{Q}}\left[\widetilde{L}_{S}(f)\right]$. Given a prior $P$ over $\mathcal{F}$ and $\delta\in(0,1)$, with probability at least $1-\delta$ over training i.i.d. samples $S\sim\mathcal{S}^{n}$, for all $Q$ over $\mathcal{F}$, we have:
$\frac{1}{B}\widetilde{L}(Q)\leq\mathrm{kl}^{-1}\left(\frac{1}{B}\widetilde{L}_{S}(Q),\frac{\mathrm{KL}\left(Q\|P\right)+\log\left(\frac{2\sqrt{n}}{\delta}\right)}{n}\right),$
where $B:=\frac{1}{\tau}+\log\left(me^{\frac{1}{\tau}}+\varepsilon\right)$.
@@
Proof of lemma 7.
To prove lemma 7, we first need the following useful lemma.
@@
Let $\alpha\in(0,1)$ and $\delta\in(0,1)$. If for any event $A(f,x,X)$ with $f\sim Q$, we have
${\mathbb{P}}_{X}\left[A(f,x,X)\mid f,x\right]\leq\delta$, then, with probability at least $1-\delta^{\alpha}$ over dataset $S\sim\mathcal{S}^{n}$,
@@
Proof of lemma 8.
For ease of notation, define $U=\mathbb{I}_{\{A(f,x,X)\}}$ with $f\sim Q$. Assume for any augmented sample $x$, ${\mathbb{P}}_{X}\left[A(f,x,X)\mid f,x\right]\leq\delta$. Then, we derive:
@@
$\leq\frac{1}{\delta^{1-\alpha}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{f,x_{i}}\left[\delta\right]=\delta^{\alpha},$
where the first inequality uses the Markov’s inequality. ∎
Define: $A(f,x,X):=\{S_{m-1}(x)-\mu(x)\leq-\varepsilon\}$ Similarly to the inequality from lemma 5, we can derive, for an augmented sample $x$, $\mathbb{P}\left(S(x,X)-\mu(x)\leq-\varepsilon\mid x\right)\leq\delta$. Next, using similar arguments to the proof of lemma lemma 6, we derive an upper bound on ~$\tilde{\ell}(x,x^{+})$:
~$\tilde{\ell}(x,x^{+})\leq B\frac{\mathbb{I}_{\{S(x,X)+\varepsilon\leq\mu(x)\}}+\mathbb{I}_{\{S(x^{+},X^{\prime})+\varepsilon\leq\mu(x^{+})\}}}{2}+\ell^{\prime}(x,x^{+},X^{-}).$
@@
Invoking lemma 8, we obtain: with probability at least $1-\delta^{\alpha}$ over dataset $S$:
@@
Let $\alpha\in(0,1)$. We will now combine all previously established lemmas to finish the proof of theorem 4. Given that $B_{\ell}\leq B$ and for $\delta_{1}\in(0,1)$, lemma 6 can be rewritten as $L(Q)\leq\widetilde{L}(Q)+B\delta_{1}$. Next, combining this inequality with corollary 1, we obtain for $\delta_{2}\in(0,1)$: with probability at least $1-\delta_{2}$ over training i.i.d. samples $S\sim\mathcal{S}^{n}$, for all $Q$ over $\mathcal{F}$,
@@
Using a union bound argument, and given the property of $\mathrm{kl}^{-1}$ as a monotonically increasing function of its first argument when fixing the second argument [35], we can now combine the previous bounds with lemma 7: with probability at least $1-\delta_{2}-\delta_{1}^{\alpha}$ over dataset $S$, for all $Q$ over $\mathcal{F}$,
@@ 3.2 Bound on the Downstream Classification Loss
Consider the cross-entropy loss $L_{\mathrm{CE}}(f,W)$ in (3). For all $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$, the following inequality holds:
@@
with $\beta(f,\sigma)=\frac{\sigma}{\tau}+L(f)+\Delta$ for $\Delta=\log\left(\frac{C}{m-1}\operatorname{cosh}^{2}\left(\frac{1}{\tau}\right)\right)$. Furthermore, $\alpha$ is given by
$\alpha=\log(C)+\min\{0,\log(\operatorname{cosh}^{2}(1))-\tau\Delta\}$ and $\sigma={\mathbb{E}}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]\leq 2$ represents the intra-class feature deviation.
@@
Practical implementations often consider a deterministic or learnable projection of the representation $f(x)$. After contrastive training, the projection head is removed, and only the backbone features are used for downstream tasks. Notably, theorem 5 can be extended to include a simple projection head of the form $(I_{k},0)$. Let $f(x)\in\mathbb{R}^{d}$ denote the backbone features and $f_{1}(x)\in\mathbb{R}^{k}$ the projection head output, where typically $d\geq k$. We can generalize theorem 5 by considering the contrastive loss on the projected features $L(f_{1})$ instead of $L(f)$ and by modifying the supervised loss $L_{\mathrm{CE}}(f,W)$ as follows:
@@
where $f_{2}(x)\in\mathbb{R}^{d-k}$ is defined such that $f(x)=\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\end{bmatrix}$.
@@
Note that theorem 5 can be applied to the original SimCLR loss by replacing $m-1$ with $2(m-1)$.
The rest of the subsection proves theorem 5. First, observe that the SimCLR population loss defined in eq. 2 can be rewritten as $L(f)=\underset{(x_{i},x_{i}^{+})_{i=1}^{m}\sim\mathcal{S}^{m}}{\mathbb{E}}\left[\frac{1}{m}\sum_{i=1}^{m}\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},X_{i})\right]$ where $X_{i}=\bigcup_{j\neq i}\{x_{j}\}$ denotes the set of negative samples. The first part of the theorem builds on the proof from Bao et al. [4], but instead of their simple framework, we enhance the applicability of the proof for the SimCLR framework, where the contrastive loss is computed over a batch and includes temperature scaling. Additionally, to address the fact that the positive pair is generated through data augmentation, we incorporate an argument from Wang et al.’s proof [43]. The following lemma presents the first part of the theorem.
@@
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\frac{\sigma}{\tau}+L(f)+\log\left(\frac{C}{m-1}\operatorname{cosh}^{2}\left(\frac{1}{\tau}\right)\right).$
@@
Let $\operatorname{LSE}(\mathbf{z}):=\log\left(\sum_{j}\exp\left(z_{j}\right)\right)$ where $\mathbf{z}:=\left\{\frac{f(x)^{\top}f(x^{\prime})}{\tau}\right\}_{x^{\prime}\in X}$ We have
$L(f)$ $\overset{(a)}{\geq}-\mathbb{E}_{S\sim\mathcal{S}^{m}}\left[\frac{1}{m}\sum_{i=1}^{m}\frac{f({x_{i}})^{\top}f({x_{i}}^{+})}{\tau}\right]+\mathbb{E}_{S\sim\mathcal{S}^{m}}\left[\frac{1}{m}\sum_{i=1}^{m}\log\sum_{x^{\prime}\in X_{i}}\exp\left(\frac{f({x_{i}})^{\top}f({x^{\prime}})}{\tau}\right)\right]$
=$\overset{(b)}{=}-\mathbb{E}_{(x,x^{+})\sim\mathcal{S}}\left[\frac{f({x})^{\top}f({x}^{+})}{\tau}\right]+\frac{1}{m}\sum_{i=1}^{m}\underset{S\sim\mathcal{S}^{m}}{\mathbb{E}}\left[\operatorname{LSE}(\mathbf{z})\right]$
$\overset{(c)}{\geq}-\frac{\mathbb{E}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{X}\left[\operatorname{LSE}(\mathbf{z})\right]$
=$\overset{(d)}{=}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{x^{\prime},y^{\prime}\}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z})\right]$
$\overset{(e)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{y^{\prime}\}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$
where (a) relies on $\exp\left(f({x})^{\top}f({x}^{+})\right)\geq 0$, (b) follows from rewriting the expectations and using the definition of the $\operatorname{LSE}$ function (c) uses the argument from Wang et al. : using the Cauchy-Schwarz inequality, assuming $y$ is the label of the positive pair $(x,x^{+})$, we can show for any $\mu_{y}$, $\left|f(x)^{\top}\frac{f(x^{+})-\mu_{y}}{\left\|f(x^{+})-\mu_{y}\right\|}\right|\leq 1$ [43], (d) uses the definition of the intra-class feature deviation $\sigma:=\mathbb{E}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]$, (e) uses the Jensen’s inequality and the convexity of the $\operatorname{LSE}$ function where we define $\mathbf{z}^{\mu}:=\left\{\frac{f(x)^{\top}\mu_{y^{\prime}}}{\tau}\right\}_{y^{\prime}}$
Next, we need to lower-bound the quantity $\mathbb{E}_{\begin{subarray}{c}\{y^{\prime}\}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$. For this, we will use the following lemma.

### v4

@@ 3.1.2 Second PAC-Bayes certificate
While theorem˜3 uses the bounded difference assumption, an alternative approach would be to replace $\sum_{x^{\prime}\in X}\operatorname{sim}(x_{i},x^{\prime})$ in (1) by its expected value, conditioned on $x_{i}$, thereby making the terms in (1) independent. To implement this in the PAC-Bayes framework, we define an $\varepsilon$-modified SimCLR empirical loss $L_{S}^{\prime}(Q)=\int_{\mathcal{W}}L_{S}^{\prime}(f_{w})\,Q(dw)$ by replacing $\ell_{\text{cont}}$ in (1) with an upper bound $\ell^{\prime}_{\text{cont}}$ as
@@
where $B:=\frac{1}{\tau}+\log\left(me^{\frac{1}{\tau}}\right)+\varepsilon$ for $\varepsilon=\left(e^{\frac{1}{\tau}}-e^{-\frac{1}{\tau}}\right)\sqrt{\frac{m-1}{2\alpha}\log\frac{2}{\delta}}$ .
@@
Observe that Theorem˜4 still holds for the original SimCLR loss by replacing $\varepsilon$ with $2\varepsilon$. This result follows from grouping terms involving augmented views of the same sample into one variable, $\operatorname{sim}(x,x^{\prime})+\operatorname{sim}(x,{x^{\prime}}^{+})$, before applying the concentration bound. This ensures that the variables are independent while multiplying the bound by $2$.
The remainder of this subsection is dedicated to the proof of theorem˜4. First, we apply a Hoeffding’s inequality to the term involving the negative samples, as stated below.
@@
Let $f\sim Q$ and recall the notation $S(x,X):=\sum\limits_{x^{\prime}\in X}\operatorname{sim}(x,x^{\prime})$ . For all $\delta\in(0,1)$, with $\varepsilon=\left(e^{\frac{1}{\tau}}-e^{-\frac{1}{\tau}}\right)\sqrt{\frac{m-1}{2}\log\frac{1}{\delta}}$, the following concentration bound holds: $\mathbb{P}\left(S(x,X)-\mathbb{E}[S(x,X)]\geq\varepsilon\mid x\right)\leq\delta.$
@@
Recall the SimCLR population loss $L(f)$ defined in eq.˜2. Define the intermediate loss $\widetilde{L}(f)$ which is a mean of i.i.d. terms obtained by replacing $\ell_{\text{cont}}$ with the loss $\tilde{\ell}_{\text{cont}}$ given by $\tilde{\ell}_{\text{cont}}(x,x^{+}):=-\log\frac{\operatorname{sim}(x,x^{+})}{\operatorname{sim}(x,x^{+})+\mathbb{E}\left[S(x,X)\mid x\right]+\varepsilon}$ for a suitable $\varepsilon>0$. We obtain the following
@@
Using lemma˜5, the following lemma shows that $L(f)$ can be upper-bounded by $\widetilde{L}(f)$.
@@
Let $f\sim Q$ and $\widetilde{L}(f)$ defined in eq.˜6. Let $\delta\in(0,1)$ and $B_{\ell}=\frac{2}{\tau}+\log m$. Then, we have
@@
Recall that $\ell_{\operatorname{cont}}(x,x^{+},X^{-})$ can be written as an average of two terms that play a symmetric role. Hence, we will first upper-bound $\ell_{\operatorname{cont}}(x,x^{+},X)$ where $X=\bigcup_{x^{\prime}\neq x}\{x^{\prime}\}$. Observe that $\ell_{\operatorname{cont}}(x,x^{+},X)$ can be split into two terms using $\varepsilon$ defined in lemma˜5:
$\ell_{\operatorname{cont}}(x,x^{+},X)\leq\tilde{\ell}_{\text{cont}}(x,x^{+})\mathbb{I}_{\{S(x,X)<\mu(x)+\varepsilon\}}+\ell_{\operatorname{cont}}(x,x^{+},X)\mathbb{I}_{\{S(x,X)\geq\mu(x)+\varepsilon\}},$
where $\mu(x)=\mathbb{E}\left[S(x,X)\mid x\right]$. Noting that $\ell_{\operatorname{cont}}(x,x^{+},X)\leq B_{\ell}$ and $\mathbb{I}_{\{\cdot\}}\leq 1$, we obtain $\ell_{\operatorname{cont}}(x,x^{+},X)\leq\tilde{\ell}_{\text{cont}}(x,x^{+})+B_{\ell}\mathbb{I}_{\{S(x,X)\geq\mu(x)+\varepsilon\}}$. We derive:
@@
where $X^{\prime}=\bigcup_{x^{\prime}\neq x}\{{x^{\prime}}^{+}\}$. Using lemma˜5 and the expression of $L(f)$ combined with eq.˜8,
$L(f)\leq\underset{(x_{i},x_{i}^{+})_{i=1}^{m}\sim\mathcal{S}^{m}}{\mathbb{E}}\left[\frac{1}{m}\sum_{i=1}^{m}\tilde{\ell}(x_{i},x_{i}^{+})\right]+B_{\ell}\delta\,$, (9)
@@
Since $\widetilde{L}_{S}(f)$ is an average of i.i.d. terms, the following result is a direct application of the PAC-Bayes-kl bound (theorem˜1) after rescaling of the loss function to the interval $[0,1]$.
@@
For $S\sim\mathcal{S}^{n}$, let $\widetilde{L}(f)$ and $\widetilde{L}_{S}(f)$ be defined as in eq.˜6 and eq.˜10 respectively, and define the extended losses: $\widetilde{L}(Q)=\mathbb{E}_{f\sim\mathcal{Q}}\left[\widetilde{L}(f)\right]$ and $\widetilde{L}_{S}(Q)=\mathbb{E}_{f\sim\mathcal{Q}}\left[\widetilde{L}_{S}(f)\right]$. Let $B:=\frac{1}{\tau}+\log(me^{\frac{1}{\tau}}+\varepsilon)$. Given a prior $P$ over $\mathcal{F}$ and $\delta\in(0,1)$, with probability at least $1-\delta$ over i.i.d. samples $S\sim\mathcal{S}^{n}$, for all $Q$ over $\mathcal{F}$,
$\frac{1}{B}\widetilde{L}(Q)\leq\mathrm{kl}^{-1}\left(\frac{1}{B}\widetilde{L}_{S}(Q),\frac{\mathrm{KL}\left(Q\|P\right)+\log\left(\frac{2\sqrt{n}}{\delta}\right)}{n}\right).$
@@
Proof of lemma˜7.
To prove lemma˜7, we first need the following useful lemma.
@@
Let $\alpha\in(0,1)$ and $\delta\in(0,1)$. If for any event $A(f,x,X)$ with $f\sim Q$, we have ${\mathbb{P}}_{X}\left[A(f,x,X)\mid f,x\right]\leq\delta$, then, with probability at least $1-\delta^{\alpha}$ over dataset $S\sim\mathcal{S}^{n}$,
@@
Proof of lemma˜8.
For ease of notation, define $U=\mathbb{I}_{\{A(f,x,X)\}}$ with $f\sim Q$. Assume for any augmented sample $x$, ${\mathbb{P}}_{X}\left[A(f,x,X)\mid f,x\right]\leq\delta$. Then, using Markov’s inequality, we derive:
@@
$\leq\frac{1}{\delta^{1-\alpha}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{f,x_{i}}\left[\delta\right]=\delta^{\alpha}.$
Define $A(f,x,X):=\{S(x,X)-\mu(x)\leq-\varepsilon\}$. Similarly to the inequality from lemma˜5, we can derive, for an augmented sample $x$, $\mathbb{P}\left(S(x,X)-\mu(x)\leq-\varepsilon\mid x\right)\leq\delta$. Next, using similar arguments to the proof of lemma lemma˜6, we derive an upper bound on $\tilde{\ell}(x,x^{+})$:
$\tilde{\ell}(x,x^{+})\leq B\frac{\mathbb{I}_{\{S(x,X)+\varepsilon\leq\mu(x)\}}+\mathbb{I}_{\{S(x^{+},X^{\prime})+\varepsilon\leq\mu(x^{+})\}}}{2}+\ell^{\prime}(x,x^{+},X^{-}).$
@@
Invoking lemma˜8, we obtain: with probability at least $1-\delta^{\alpha}$ over dataset $S$:
@@
Let $\alpha\in(0,1)$. We will now combine all previously established lemmas to finish the proof of theorem˜4. Given that $B_{\ell}\leq B$ and for $\delta_{1}\in(0,1)$, lemma˜6 can be rewritten as $L(Q)\leq\widetilde{L}(Q)+B\delta_{1}$. Next, combining this inequality with corollary˜1, we obtain for $\delta_{2}\in(0,1)$: with probability at least $1-\delta_{2}$ over training i.i.d. samples $S\sim\mathcal{S}^{n}$, for all $Q$ over $\mathcal{F}$,
@@
Using a union bound argument, and given the property of $\mathrm{kl}^{-1}$ as a monotonically increasing function of its first argument when fixing the second argument [41], we can now combine the previous bounds with lemma˜7: with probability at least $1-\delta_{2}-\delta_{1}^{\alpha}$ over dataset $S$, for all $Q$ over $\mathcal{F}$,
@@ 3.2 Bound on the Downstream Classification Loss
Consider the cross-entropy loss $L_{\mathrm{CE}}(f,W)$ in (3) and denote the class distribution by $\boldsymbol{\pi}:=[\mathbb{P}(Y=c)]_{c\in\mathcal{C}}$. For all $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$, the following inequality holds:
@@
with $\beta(f,\sigma)=\frac{\sigma}{\tau}+L(f)+\Delta$ for $\Delta=\log\left(\frac{C^{2}\pi^{*}}{m-1}\operatorname{cosh}^{2}\left(\frac{1}{\tau}\right)\right)$ where $\pi^{*}:=\max_{c\in\mathcal{C}}\pi(c)$.
Furthermore, $\alpha$ is given by $\alpha=\log(C)+\min\{0,\log(C\pi^{*}\operatorname{cosh}^{2}(1))-\tau\Delta\}$ and $\sigma={\mathbb{E}}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]\leq 2$ represents the intra-class feature deviation.
@@
Practical implementations often consider a deterministic or learnable projection of the representation $f(x)$. After contrastive training, the projection head is removed, and only the backbone features are used for downstream tasks. Notably, theorem˜5 can be extended to include a simple projection head of the form $(I_{k},0)$. Let $f(x)\in\mathbb{R}^{d}$ denote the backbone features and $f_{1}(x)\in\mathbb{R}^{k}$ the projection head output, where typically $d\geq k$. We can generalize theorem˜5 by considering the contrastive loss on the projected features $L(f_{1})$ instead of $L(f)$ and by modifying the supervised loss $L_{\mathrm{CE}}(f,W)$ as follows:
@@
where $f_{2}(x)\in\mathbb{R}^{d-k}$ is defined such that $f(x)=\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\end{bmatrix}$ and $W=\begin{bmatrix}W^{(1)}\\ W^{(2)}\end{bmatrix}$.
@@
Note that theorem˜5 can be applied to the original SimCLR loss by replacing m1m-1 with 2(m1)2(m-1). When the class distribution is uniform, $\pi^{*}$ simplifies to $1/C$.
The rest of the subsection proves theorem˜5. First, observe that the SimCLR population loss defined in eq.˜2 can be rewritten as $L(f)={\mathbb{E}}_{(x_{i},x_{i}^{+})_{i=1}^{m}\sim\mathcal{S}^{m}}\left[\frac{1}{m}\sum_{i=1}^{m}\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},X_{i})\right]$ where $X_{i}=\bigcup_{j\neq i}\{x_{j}\}$ denotes the set of negative samples. The first part of the theorem builds on the proof from Bao et al. [3], but instead of their simple framework, we enhance the applicability of the proof for the SimCLR framework, where the contrastive loss is computed over a batch and includes temperature scaling. Additionally, to address the fact that the positive pair is generated through data augmentation, we incorporate an argument from Wang et al.’s proof [52]. The following lemma presents the first part of the theorem.
@@
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\frac{\sigma}{\tau}+L(f)+\Delta.$
@@
Let $\operatorname{LSE}(\mathbf{z}):=\log\left(\sum_{j}\exp\left(z_{j}\right)\right)$ where $\mathbf{z}:=\left\{{f(x)^{\top}f(x^{\prime})/\tau}\right\}_{x^{\prime}\in X}$. For convenience, we define $X$ as the set of negative samples corresponding to the anchor sample $x$, with the total number of negative samples denoted by K=m1K=m-1. The elements of $X$ are indexed by $i=1,\ldots,K$. Denoting the class-conditional distribution by $\mathcal{D}_{c}:=\mathbb{P}(X\mid$ $Y=c)$ for each $c\in\mathcal{C}$, we define $\mu_{c}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{c}}[f(x)]$ and note that $\|\mu_{c}\|\leq 1$. We have
$L(f)$ $\overset{(a)}{\geq}-\underset{{S\sim\mathcal{S}^{m}}}{\mathbb{E}}\left[\frac{1}{m}\sum_{i=1}^{m}\frac{f({x_{i}})^{\top}f({x_{i}}^{+})}{\tau}\right]+\underset{{S\sim\mathcal{S}^{m}}}{\mathbb{E}}\left[\frac{1}{m}\sum_{i=1}^{m}\log\sum_{x^{\prime}\in X_{i}}\exp\left(\frac{f({x_{i}})^{\top}f({x^{\prime}})}{\tau}\right)\right]$ (11)
$\overset{(b)}{=}-\mathbb{E}_{(x,x^{+})\sim\mathcal{S}}\left[\frac{f({x})^{\top}f({x}^{+})}{\tau}\right]+\frac{1}{m}\sum_{i=1}^{m}\underset{S\sim\mathcal{S}^{m}}{\mathbb{E}}\left[\operatorname{LSE}(\mathbf{z})\right]$ (12)
$\overset{(c)}{\geq}-\frac{\mathbb{E}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{X}\left[\operatorname{LSE}(\mathbf{z})\right]$ (13)
$\overset{(d)}{=}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{(x_{i}^{\prime},y_{i}^{\prime})\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z})\right]$ (14)
$\overset{(e)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{y_{i}^{\prime}\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$ (15)
where (a) relies on $\exp\left(f({x})^{\top}f({x}^{+})\right)\geq 0$, (b) follows from rewriting the expectations and using the definition of the $\operatorname{LSE}$ function (c) uses the argument from Wang et al. : using the Cauchy-Schwarz inequality, assuming $y$ is the label of the positive pair $(x,x^{+})$, we can show for any $\mu_{y}$, $\left|f(x)^{\top}\frac{f(x^{+})-\mu_{y}}{\left\|f(x^{+})-\mu_{y}\right\|}\right|\leq 1$ [52], (d) uses the definition of the intra-class feature deviation $\sigma:=\mathbb{E}_{(x,y)}\left[\|f(x)-\mu_{y}\|_{2}\right]$, (e) uses the Jensen’s inequality and the convexity of the $\operatorname{LSE}$ function where we define $\mathbf{z}^{\mu}:=\left\{{f(x)^{\top}\mu_{y^{\prime}}/\tau}\right\}_{y^{\prime}\in Y}$ where we denote $Y=\begin{subarray}{c}\{y_{i}^{\prime}\}_{i=1}^{K}\end{subarray}$.
Next, we need to lower-bound the quantity $\mathbb{E}_{Y}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$. We will use the following lemma.